# Tunable Molecular Interactions Near an Atomic Feshbach Resonance: Stability and Collapse of a Molecular Bose-Einstein Condensate

### v1

@@ Abstract
where $\{\Psi_{\sigma 0},\chi_{\sigma\mathbf{k}}\}$ are the variational parameters and $\mathcal{N}$ is the normalization factor. In the exponent, the $\mathbf{k}-$sum is over half of $\mathbf{k}-$space, and the prime in the $\mathbf{k}$-summation implies the origin $\mathbf{k}=0$ is excluded. The vacuum $|0\rangle$ satisfies $a_{\sigma\mathbf{k}}|0\rangle=0$ by all annihilation operators $a_{\sigma\mathbf{k}}$. In the spirit of generalized Bogoliubov theory, this variational wavefunction includes only pair-wise correlations between atoms or between molecules in the exponent, which is sufficient for our focus on the detuning regime that is not too close to the resonance Wang et al. (2025).
@@
We now numerically solve the five equations. Figures 3(a,b) plot the calculated chemical potentials near resonance at two different densities for both the wide and narrow resonances. This provides useful insight into the anomalous negative sign of the compressibility and the related condensate instability. From both figures we observe that the chemical potential $\mu$, which represents the average energy per atom, falls below the two energy levels corresponding to the bare atomic continuum threshold on the atomic side and the two-body molecular energy$/2$ on the molecular side. Furthermore, as the density $n$ decreases, in the near-resonance regime the chemical potential $\mu$ approaches the two energy levels from below. Consequently, the inverse compressibility $\kappa^{-1}=d\mu/dn$ must be negative near and on both sides of the resonance  [^7] In contrast to the near-resonance regime in Figs. 3(a,b), the chemical potential $\mu$ rises above one of the two energy levels when the detuning is far off the resonance. This is true for both the wide- and narrow-resonance cases (see Ref. sup for additional details). It implies that, in these far-off-resonance regimes, the chemical potential must approach these levels from above as the density vanishes $n\rightarrow 0$, consistent with the fact that the compressibility is positive in those detuning regimes. . The numerical results which were shown previously in Figs. 2(a,b) support this analysis; there we directly evaluated the derivative $d\mu/dn$ numerically, from which we constructed the ground-state stability phase diagrams.
Figure 3: Chemical potential, inverse compressibility and molecular scattering lengths. (a,b) Behavior of the chemical potential $\mu$ (blue solid and red dashed lines) versus detuning $\bar{\nu}$ at two different densities, $n$ and $n/2$ with $na_{\mathrm{bg}}^{3}=1.68\times 10^{-5}$ the same density value used in Figs. 2(c,d). “Atom continuum” threshold corresponds to $\mu=0$, and “$-E_{b}$” in (a) is the two-body but dressed molecular energy, which is replaced by its bare value $\bar{\nu}$ in (b). (c,d) Comparison of the corresponding inverse compressibility (blue solid lines) for the same density $na_{\mathrm{bg}}^{3}=1.68\times 10^{-5}$ with the atomic scattering length $a_{s}$ (red dashed lines) and the many-body molecular scattering length $a_{\mathrm{mm}}^{\mathrm{eff}}$ (brown dashed lines with squares), which is defined in Eq. (9). Both $a_{\mathrm{mm}}^{\mathrm{eff}}$ and $a_{s}$ are plotted in units of $a_{\mathrm{bg}}$.
@@
with $\text{with }\quad g_{2}^{\mathrm{eff}}$ $=\frac{(\mathcal{Z}^{\mathrm{eff}})^{2}}{2}\frac{\partial^{4}\Omega[\Psi_{10},\Psi_{20},\Delta_{1},\Delta_{2},\mu]/V}{\partial\Psi_{20}^{2}\partial(\Psi_{20}^{*})^{2}}.$ (9b)
Here, $\mathcal{Z}^{\mathrm{eff}}$ is the fraction of the dressed-molecular condensate in the closed channel, defined by $\mathcal{Z}^{\mathrm{eff}}=|\Psi_{20}|^{2}/(|\Psi_{20}|^{2}+n_{1}/2)$  [^8] The factor $(\mathcal{Z}^{\mathrm{eff}})^{2}$ in Eq. (9b) is needed to reconcile two facts: (1) the many-body scattering length $a_{\mathrm{mm}}^{\mathrm{eff}}$, (as well as the two-body one $a_{\mathrm{mm}}$), is defined for a dressed molecule sup , which is an admixture of closed-channel molecules with open-channel atom pairs, and (2) the derivative in Eq. (9b) is taken with respect to the bare closed-channel molecular condensate amplitude $\Psi_{20}$. Although this factor is essentially equal to $1$ for a narrow resonance, it becomes important for wide resonances, particularly in deriving Eq. (2sup in the zero density limit. . The quartic derivative in Eq. (9b) can be carried out approximately sup  [^9] When performing derivatives in Eq. (9b) we set $\{\Psi_{10},\Delta_{1},\Delta_{2}\}$ at their saddle-point values. They are viewed as functions of $\{\Psi_{20},\mu\}$ only. In arriving at Eq. (10) we have also neglected terms that are higher order in the ratios $\{g_{1}n/E_{1\mathbf{k}=0},g_{2}n/E_{1\mathbf{k}=0},\alpha\sqrt{n}/E_{1\mathbf{k}=0}\}$. , leading to
$\frac{g_{2}^{\mathrm{eff}}}{(\mathcal{Z}^{\mathrm{eff}})^{2}}\approx\widetilde{\bar{g}}_{2}-\widetilde{\bar{\alpha}}^{4}\big{[}\frac{1}{V}\sum_{\mathbf{k}}\frac{1}{2E_{1\mathbf{k}}^{3}}-2g_{1}\big{(}\frac{1}{V}\sum_{\mathbf{k}}\frac{\widetilde{\epsilon}_{1\mathbf{k}}}{2E_{1\mathbf{k}}^{3}}\big{)}^{2}\big{]},$ (10)
where $E_{1\mathbf{k}}=\sqrt{\widetilde{\epsilon}_{1\mathbf{k}}^{2}-|\widetilde{\Delta}_{1}|^{2}}$ is the atomic Bogoliubov quasiparticle energy, with $\widetilde{\epsilon}_{1\mathbf{k}}=h_{1\mathbf{k}}+2g_{1}(|\Psi_{10}|^{2}+n_{1})$ and $\widetilde{\Delta}_{1}=\Delta_{1}+g_{1}\Psi_{10}^{2}-2\alpha\Psi_{20}$. In Eq. (10), $\widetilde{\bar{\alpha}}$ and $\widetilde{\bar{g}}_{2}$ are two interaction parameters, related to the Feshbach coupling $\bar{\alpha}$ and molecule-molecule interaction $\bar{g}_{2}$ by $\widetilde{\bar{\alpha}}=\sqrt{2}\alpha/(1+g_{1}V^{-1}\sum_{\mathbf{k}}1/2E_{1\mathbf{k}})$ and $\widetilde{\bar{g}}_{2}=g_{2}/(1+g_{2}V^{-1}\sum_{\mathbf{k}}1/2E_{2\mathbf{k}})$. It is important to note that Eq. (10) is applicable as long as the detuning $\bar{\nu}$ is smaller than and not too close to its QCP value $\bar{\nu}_{c}$ as we have already set $\Psi_{10}=0$. Evaluating Eqs. (9) and (10) leads to the $a_{\mathrm{mm}}^{\mathrm{eff}}$ plots in Figs. 2(c,d).
@@
In general, the $\mathbf{k}$ integral in Eq. (10) can not be done analytically. However, for the narrow resonance case and detuning near the QCP, Eqs. (10) and (9) can be further simplified to yield the simple analytical expression for the many-body scattering length $a_{\mathrm{mm}}^{\mathrm{eff}}$ sup presented in Eq. (4). In Figs. 3(c,d), we numerically evaluate $a_{\mathrm{mm}}^{\mathrm{eff}}$ for a generic detuning in both the narrow and wide resonance cases, and compare it with the numerically calculated inverse compressibility $\kappa^{-1}$   [^10] At large negative detuning $\bar{\nu}$, the many-body scattering length $a_{\mathrm{mm}}^{\mathrm{eff}}$ approaches its two-body background counterpart $a_{\mathrm{mm}}^{\mathrm{bg}}$.. Notably, the two, $a_{\mathrm{mm}}^{\mathrm{eff}}$ and $\kappa^{-1}$, show rather precise agreement, for both the narrow and wide resonances at detunings away from the immediate vicinity of the QCP  [^11] We note that the more extended unstable regime in Fig. 2(a) for a wide resonance is associated with the presence of a sizable number of atom Cooper pairs sup , which is also a consequence of the larger Feshbach coupling strength $\bar{\alpha}$. .
All of this allows us to understand why the compressibility in Fig. 2(a,b) behaves so differently when comparing the behavior of wide and narrow resonances. For a very narrow resonance, because of the small Feshbach coupling $\bar{\alpha}$, the factor $\widetilde{\bar{\alpha}}^{4}$ in Eq. (10), which is proportional to $\bar{\alpha}^{4}$, is very small. Consequently the effective scattering length $a_{\mathrm{mm}}^{\mathrm{eff}}$ becomes negative only when the atomic Bogoliubov quasiparticle energy gap, (contained in $E_{1\mathbf{k}=0}$), is sufficiently small. This insures that the detuning at which $a_{\mathrm{mm}}^{\mathrm{eff}}$ changes sign is sufficiently close to the QCP. By contrast, for a wide resonance this occurs when the gap $E_{1\mathbf{k}=0}$ is still large, which corresponds to a detuning well away from the QCP.
@@
where $r_{*}$ was defined in Eq. (6) and $k_{b}=\sqrt{mE_{b}}/\hbar$ is the detuning-dependent momentum corresponding to the molecular binding energy $E_{b}$. The three dimensionless and positive coefficients $\{A_{1},A_{2},A_{3}\}$ contain sub-leading dependences on $1/k_{b}$ sup , and their expressions are given in the Supplemental Materials sup . To arrive at Eqs. (2) and (3), we retain the background contribution $a_{\mathrm{mm}}^{\mathrm{bg}}$ in Eq. (11) along with the $A_{1}$ and $A_{2}$ terms, which are dominant for the wide and narrow resonances, respectively, near the boundary detuning $\bar{\nu}_{c,-}$ in Fig. 2.
Conclusions — In this Letter we have presented an in-depth study of the stability of molecular condensates in the context of Feshbach resonances of bosonic atoms. This has been a longstanding challenge for the field. By numerically calculating the compressibility we have not only constructed ground-state phase diagrams, but also derived concrete expressions for the two-body molecular scattering length $a_{\mathrm{mm}}$ and its many-body analogue $a_{\mathrm{mm}}^{\mathrm{eff}}$, given in Eqs. (2), (3), (4), and (11). These formulae, which through sign changes, allow one to infer the instability regions of molecular condensates, can and should be directly tested in future experiments. The concrete expressions for the molecular scattering lengths we derived make it possible to address a number of issues which were previously inaccessible; these include, among others, the equilibrium equation of state in the dilute molecular gas regime, the collective modes, the molecule-molecule scattering induced relaxation Petrov et al. (2005); Levinsen and Gurarie (2006); Levinsen and Petrov (2011), and the formation of the tetramer bound state Chin et al. (2005); Chen et al. (2023).
Acknowledgment — We thank P. Julienne and B. Zhao for helpful discussions. Z. W. and Q. C. are supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301904). This work is also supported by the National Science Foundation under Grant No. PHY-2409612 and PHY-2103542, and by the Air Force Office of Scientific Research under award number FA9550-21-1-0447. Z. Z. acknowledges the Bloch Postdoctoral Fellowship. We also acknowledge the University of Chicago’s Research Computing Center for their support of this work.

### v2

@@ Abstract
where $\{\Psi_{\sigma 0},\chi_{\sigma\mathbf{k}}\}$ are the variational parameters and $\mathcal{N}$ is the normalization factor. In the exponent, the $\mathbf{k}$-sum is over half of $\mathbf{k}$-space, and the prime in the $\mathbf{k}$-summation implies the origin $\mathbf{k}=0$ is excluded. The vacuum $|0\rangle$ satisfies $a_{\sigma\mathbf{k}}|0\rangle=0$ for all annihilation operators $a_{\sigma\mathbf{k}}$. In the spirit of generalized Bogoliubov theory, this variational wavefunction includes only pair-wise correlations between atoms or between molecules in the exponent. We emphasize that this approximation is adequate for our focus on the detuning regime that is not too close to the resonanceWang et al. (2025). In Ref.sup , we provide concrete estimates showing that neglecting higher-order correlations results in less than $2\%$ uncertainty in the phase boundaries of the phase diagram in Fig. 2, as well as in the scattering length formulas given in Eqs. (2) –(4).
@@
We now numerically solve the five equations. Figures 3(a) and 3(b) plot the calculated chemical potentials near resonance at two different densities for both the wide and narrow resonances. This provides useful insight into the anomalous negative sign of the compressibility and the related condensate instability. From both figures we observe that the chemical potential $\mu$, which represents the average energy per atom, falls below the two energy levels corresponding to the bare atomic continuum threshold on the atomic side and the two-body molecular energy/2/2 on the molecular side. Furthermore, as the density $n$ decreases, in the near-resonance regime the chemical potential $\mu$ approaches the two energy levels from below. Consequently, the inverse compressibility $\kappa^{-1}=d\mu/dn$ must be negative near and on both sides of the resonance  [^7] In contrast to the near-resonance regime in Figs. 3(a,b), the chemical potential $\mu$ rises above one of the two energy levels when the detuning is far off the resonance. This is true for both the wide- and narrow-resonance cases (see Ref. sup for additional details). It implies that, in these far-off-resonance regimes, the chemical potential must approach these levels from above as the density vanishes ($n\rightarrow 0$), consistent with the fact that the compressibility is positive in those detuning regimes. . The numerical results that were shown previously in Figs. 2(a) and 2(b) support this analysis; there we directly evaluated the derivative $d\mu/dn$ numerically, from which we constructed the ground-state stability phase diagrams.
Figure 3: Chemical potential, inverse compressibility and molecular scattering lengths. (a,b) Chemical potential $\mu$ (blue solid and red dashed lines) vs detuning $\bar{\nu}$ at densities $n$ and $n/2$, with $na_{\mathrm{bg}}^{3}=1.68\times 10^{-5}$ [same as Figs. 2(c) and 2(d)]. The “atom continuum” threshold is at $\mu=0$. In panel (a), $E_{b}$” represents the dressed molecular energy; in panel (b), it is replaced by its bare value, $\bar{\nu}$. (c), (d) Comparison of the inverse compressibility (blue solid lines) at $na_{\mathrm{bg}}^{3}=1.68\times 10^{-5}$ with the atomic scattering length $a_{s}$ (red dashed lines) and the many-body molecular scattering length $a_{\mathrm{mm}}^{\mathrm{eff}}$ (brown dashed lines with squares), defined in Eq. (9). Both $a_{\mathrm{mm}}^{\mathrm{eff}}$ and $a_{s}$ are plotted in units of $a_{\mathrm{bg}}$.
@@
$\text{with }\quad g_{2}^{\mathrm{eff}}$ $=\frac{(\mathcal{Z}^{\mathrm{eff}})^{2}}{2}\frac{\partial^{4}\Omega[\Psi_{10},\Psi_{20},\Delta_{1},\Delta_{2},\mu]/V}{\partial\Psi_{20}^{2}\partial(\Psi_{20}^{*})^{2}}.$ (9b)
Here, $\mathcal{Z}^{\mathrm{eff}}$ is the fraction of the dressed-molecular condensate in the closed channel, defined by $\mathcal{Z}^{\mathrm{eff}}=|\Psi_{20}|^{2}/(|\Psi_{20}|^{2}+n_{1}/2)$  [^8] The factor $(\mathcal{Z}^{\mathrm{eff}})^{2}$ in Eq. (9b) is needed to reconcile two facts: (1) the many-body scattering length $a_{\mathrm{mm}}^{\mathrm{eff}}$, as well as the two-body one $a_{\mathrm{mm}}$, is defined for a dressed molecule sup , which is an admixture of closed-channel molecules with open-channel atom pairs, and (2) the derivative in Eq. (9b) is taken with respect to the bare closed-channel molecular condensate amplitude $\Psi_{20}$. Although this factor is essentially equal to $1$ for a narrow resonance, it becomes important for wide resonances, particularly in deriving Eq. (2sup in the zero density limit. . The quartic derivative in Eq. (9b) can be carried out approximately sup  [^9] When performing derivatives in Eq. (9b) we set $\{\Psi_{10},\Delta_{1},\Delta_{2}\}$ at their saddle-point values. They are viewed as functions of $\{\Psi_{20},\mu\}$ only. In arriving at Eq. (10) we have also neglected terms that are higher order in the ratios $\{g_{1}n/E_{1\mathbf{k}=0},g_{2}n/E_{1\mathbf{k}=0},\alpha\sqrt{n}/E_{1\mathbf{k}=0}\}$. , leading to
$\frac{g_{2}^{\mathrm{eff}}}{(\mathcal{Z}^{\mathrm{eff}})^{2}}\approx\widetilde{\bar{g}}_{2}-\widetilde{\bar{\alpha}}^{4}\bigg[\frac{1}{V}\sum_{\mathbf{k}}\frac{1}{2E_{1\mathbf{k}}^{3}}-2g_{1}\bigg(\frac{1}{V}\sum_{\mathbf{k}}\frac{\widetilde{\epsilon}_{1\mathbf{k}}}{2E_{1\mathbf{k}}^{3}}\bigg)^{2}\bigg],$ (10)
where $E_{1\mathbf{k}}=\sqrt{\widetilde{\epsilon}_{1\mathbf{k}}^{2}-|\widetilde{\Delta}_{1}|^{2}}$ is the atomic Bogoliubov quasiparticle energy, with $\widetilde{\epsilon}_{1\mathbf{k}}=h_{1\mathbf{k}}+2g_{1}(|\Psi_{10}|^{2}+n_{1})$ and $\widetilde{\Delta}_{1}=\Delta_{1}+g_{1}\Psi_{10}^{2}-2\alpha\Psi_{20}$. In Eq. (10), $\widetilde{\bar{\alpha}}$ and $\widetilde{\bar{g}}_{2}$ are two interaction parameters, related to the Feshbach coupling $\bar{\alpha}$ and molecule-molecule interaction $\bar{g}_{2}$ by $\widetilde{\bar{\alpha}}=\sqrt{2}\alpha/(1+g_{1}V^{-1}\sum_{\mathbf{k}}1/2E_{1\mathbf{k}})$ and $\widetilde{\bar{g}}_{2}=g_{2}/(1+g_{2}V^{-1}\sum_{\mathbf{k}}1/2E_{2\mathbf{k}})$. It is important to note that Eq. (10) is applicable as long as the detuning $\bar{\nu}$ is smaller than and not too close to its QCP value $\bar{\nu}_{c}$ as we have already set $\Psi_{10}=0$. Evaluating Eqs. (9) and (10) leads to the $a_{\mathrm{mm}}^{\mathrm{eff}}$ plots in Figs. 2(c) and 2(d).
@@
In general, the $\mathbf{k}$ integral in Eq. (10) can not be done analytically. However, for the narrow resonance case and detuning near the QCP, Eqs. (10) and (9) can be further simplified to yield the simple analytical expression for the many-body scattering length $a_{\mathrm{mm}}^{\mathrm{eff}}$ sup presented in Eq. (4). In Figs. 3(c) and 3(d), we numerically evaluate $a_{\mathrm{mm}}^{\mathrm{eff}}$ for a generic detuning in both the narrow and wide resonance cases, and compare it with the numerically calculated compressibility inverse $\kappa^{-1}$   [^10] At large negative detuning $\bar{\nu}$, the many-body scattering length $a_{\mathrm{mm}}^{\mathrm{eff}}$ approaches its two-body background counterpart $a_{\mathrm{mm}}^{\mathrm{bg}}$.. Notably, the two, $a_{\mathrm{mm}}^{\mathrm{eff}}$ and $\kappa^{-1}$, show rather precise agreement, for both the narrow and wide resonances at detunings away from the immediate vicinity of the QCP  [^11] We note that the more extended unstable regime in Fig. 2(a) for a wide resonance is associated with the presence of a sizable number of atom Cooper pairs sup , which is also a consequence of the larger Feshbach coupling strength $\bar{\alpha}$. .
All of this allows us to understand why the compressibility in Fig. 2(a) and 2(b) behaves so differently when comparing the behavior of wide and narrow resonances. For a very narrow resonance, because of the small Feshbach coupling $\bar{\alpha}$, the factor $\widetilde{\bar{\alpha}}^{4}$ in Eq. (10), which is proportional to $\bar{\alpha}^{4}$, is very small. Consequently the effective scattering length $a_{\mathrm{mm}}^{\mathrm{eff}}$ becomes negative only when the atomic Bogoliubov quasiparticle energy gap (contained in $E_{1\mathbf{k}=0}$) is sufficiently small. This insures that the detuning at which $a_{\mathrm{mm}}^{\mathrm{eff}}$ changes sign is sufficiently close to the QCP. In contrast, for a wide resonance this occurs when the gap $E_{1\mathbf{k}=0}$ is still large, which corresponds to a detuning well away from the QCP.
@@
where $r_{*}$ was defined in Eq. (6) and $k_{b}=\sqrt{mE_{b}}/\hbar$ is the detuning-dependent momentum corresponding to the molecular binding energy $E_{b}$. The three dimensionless and positive coefficients $\{A_{1},A_{2},A_{3}\}$ contain sub-leading dependences on $1/k_{b}$ sup , and their expressions are given in the Supplemental Material sup . To arrive at Eqs. (2) and (3), we retain the background contribution $a_{\mathrm{mm}}^{\mathrm{bg}}$ in Eq. (11) along with the $A_{1}$ and $A_{2}$ terms, which are dominant for the wide and narrow resonances, respectively, near the boundary detuning $\bar{\nu}_{c,-}$ in Fig. 2.
Conclusions– In this Letter we have conducted an in-depth study on the stability of molecular condensates near Feshbach resonances in bosonic atoms, establishing the relationship to the inter-molecular scattering lengths. In contrast with the Fermi gases, there has, thus far, been very little theoretical work on characterizing this property for those molecules comprised of bosonic atoms. Such calculations can not naturally build on past work for the Fermi systems Levinsen and Gurarie (2006) as in the Bose problem in the more immediate vicinity of resonance one has to contend with Efimov and tetramer bound states as well as other background contributions. With the caveat that, following experiment  Zhang et al. (2021, 2023), we are not too close to resonance (where instability is guaranteed) we have shown how just as Feshbach resonances allow tuning of the scattering lengths of atoms, they also modify the scattering length of molecules constituted from these atoms.
Our predictions are ripe for experimental validation. The concrete expressions for the molecular scattering lengths we derived make it possible to address a number of issues that were previously inaccessible; these include, among others, the equilibrium equation of state in the dilute molecular gas regime, the expansion dynamics, the collective modes, and the molecule-molecule scattering induced relaxation Petrov et al. (2005); Levinsen and Gurarie (2006); Levinsen and Petrov (2011); Chin et al. (2005); Chen et al. (2023).
Acknowledgment — We thank P. Julienne and B. Zhao for helpful discussions. Z. W. and Q. C. were supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301904). This work was also supported by the National Science Foundation under Grant No. PHY-2409612 and No. PHY-2103542, and by the Air Force Office of Scientific Research under Award No. FA9550-21-1-0447. Z. Z. acknowledges the Bloch Postdoctoral Fellowship. We also acknowledge the University of Chicago’s Research Computing Center for their support of this work.